Genetics problem!
Alkaptonuria is a rare recessive disorder that cau
Answer #1

Aa Aa Alkaptonuria- aa Carrier -Aa Normal Individual-AA (AM/Aa) Aa Aa AA/Aa)A/Aa)(I/2) AA or Aa A/Aa) (Aa) AA or Aa AA or Aa

1.What is the probability that individual A is heterozygous?

It will have 50% (1/2) chances of being heterozygous.

2.What is the probability that individual B is heterozygous?

The probability of B can be heterozygous can be considered from mating between II-7 and II-8. Their child III-7 will have 2/3 probability of having PKU allele and it will have ½ probability of passing it to the IV-5. It will have (2/3x1/2=1/3) chances of being heterozygous?

3.Individual B marries individual C. What is the probability that they will have a child with alkaptonuria?

If both parents (B and C) are heterozygous, they have a ¼ chance of both passing the alkaptonuria allele to their child.

Probability of getting alkaptonuria allele= p(B is A/a) x p(C is A/a) x p( both parents donate a)

= 1/3x1/4x1/4=1/48